OpenStudy (anonymous):

Help Please!!!! Ill Give Medal!! What is the solution to the equation? 2(m - 8) = 4(m + 6) ? m = ?????

OpenStudy (anonymous):

:) i have more after your done with this one :)

OpenStudy (anonymous):

2(m - 8) = 4(m + 6) The first thing we need to do is distribute. Starting with the first part. 2(m - 8) We need to multiply 2 by everything in the parenthesis. 2 * m = 2m 2 * -8 = -16 The last part: 4(m + 6) 4 * m = 4m 4 * 6 = 24 Putting this together: 2m - 16 = 4m + 24 We want to isolate m on one side, so I'm going to subtract 2m from both sides. -16 = 2m + 24 Subtract 24 from both sides. -40 = 2m Divide each side by 2. m = -20 <--- Answer. We can plug this in to see if it's correct. (: 2(m - 8) = 4(m + 6) 2 (-20 - 8) = 4(-20 + 6) 2 ( -28) = 4(-14) -56 = -56 True.

OpenStudy (anonymous):

ooOoooOOooOhhHHhh :D and thanks for explaining how you got your answer :D

OpenStudy (anonymous):

M=2

OpenStudy (anonymous):

The product of 4 and a number is 18 more than twice the number. What is the number? ????

OpenStudy (anonymous):

@BunnyBree i only have 1 more question :)

OpenStudy (anonymous):

The product of 4 and a number is 18 more than twice the number. What is the number? ????